# citations.bib

### v1

@@ H.3 Faithfulness Experiments
$\ell_{2}$ squared 0.0883 0.361 0.473
$\ell_{2}$ modified 0.101 0.414 0.54
NHANES 16384 kernel 0.00206 0.00259 0.00336
$\ell_{2}$ squared 0.0023 0.0029 0.00347
$\ell_{2}$ modified 0.0021 0.00269 0.00315
Table 5: Values for $b_{0}$ in Fig. 2 (3) Comparison: Quantile Values for Different Datasets and Sampling Methods
Dataset Samples Approximation Average MSE Time $Q_{1}$ Median MSE $Q_{3}$
MNIST 500 Matrix-Vector 0.153 0.714 0.119 0.138 0.161
Least Squares 0.0202 1.85 0.0112 0.0157 0.0309
KernelSHAP 0.0303 0.739 0.0182 0.0235 0.0304
1000 Matrix-Vector 0.0663 0.927 0.059 0.0645 0.0746
Least Squares 0.00823 1.96 0.00406 0.00718 0.0122
KernelSHAP 0.0126 0.892 0.00731 0.0101 0.0137
10000 Matrix-Vector 0.0071 5.54 0.00536 0.00712 0.0088
Least Squares 0.000659 7.71 0.000369 0.000581 0.000837
KernelSHAP 0.00112 2.94 0.000607 0.000727 0.00136
100000 Matrix-Vector 0.000802 52.4 0.000608 0.000842 0.000977
Least Squares 6.69e-05 67.9 3.68e-05 5.03e-05 8.31e-05
KernelSHAP 9.17e-05 28.4 5.46e-05 6.83e-05 9.38e-05
CIFAR10 500 Matrix-Vector 61.3 19.9 31 45.5 91.2
Least Squares 1.21 981 1.02 1.09 1.42
KernelSHAP 1.8e+06 139 9.17e+03 2.97e+04 2.75e+05
1000 Matrix-Vector 29.4 21.8 15.2 22.7 41.5
Least Squares 1.47 1.27e+03 1.06 1.22 1.82
KernelSHAP 5.18e+04 135 4.66e+03 6.07e+03 7.58e+03
10000 Matrix-Vector 2.93 56.9 1.49 2.15 4.35
Least Squares 6.6 1.21e+03 2.15 4.12 10.5
KernelSHAP 14.7 282 4.72 8.58 23.8
100000 Matrix-Vector 0.295 319 0.15 0.209 0.427
Least Squares 0.251 3.03e+03 0.0893 0.154 0.41
KernelSHAP 0.572 1.86e+03 0.211 0.355 0.917
Table 6: Performance Metrics for MNIST and CIFAR-10 Datasets Using Different Methods
@@
MNIST 500 Matrix-Vector 0.21 0.666 0.967
Least Squares 0.205 0.678 0.989
KernelSHAP 0.196 0.675 0.987
1000 Matrix-Vector 0.205 0.678 0.974
Least Squares 0.204 0.674 0.995
KernelSHAP 0.203 0.676 0.992
10000 Matrix-Vector 0.198 0.679 0.995
Least Squares 0.202 0.678 0.998
KernelSHAP 0.198 0.677 0.998
100000 Matrix-Vector 0.2 0.677 0.998
Least Squares 0.197 0.677 0.999
KernelSHAP 0.198 0.678 0.999
CIFAR10 500 Matrix-Vector 0.4 0.608 0.0568
Least Squares 0.355 0.625 0.0762
KernelSHAP 0.43 0.511 0.00512
1000 Matrix-Vector 0.406 0.61 0.104
Least Squares 0.365 0.663 0.122
KernelSHAP 0.418 0.565 0.00851
10000 Matrix-Vector 0.297 0.706 0.357
Least Squares 0.329 0.649 0.285
KernelSHAP 0.35 0.701 0.244
100000 Matrix-Vector 0.315 0.716 0.659
Least Squares 0.3 0.697 0.685
KernelSHAP 0.309 0.72 0.582
Table 7: AUC and Rank Correlation for MNIST and CIFAR10 Datasets Using Different Methods

### v2

@@ H.3 Faithfulness Experiments
$\ell_{2}$-squared 0.0883 0.361 0.473
modified $\ell_{2}$ 0.101 0.414 0.54
NHANES 50000 kernel 0.002350 0.002696 0.002915
$\ell_{2}$-squared 0.001704 0.001868 0.002084
modified $\ell_{2}$ 0.001845 0.002055 0.002203
Table 5: Values for $b_{0}$ in Fig.˜2 (3) Comparison: Quantile Values for Different Datasets and Sampling Methods (least squares estimator without replacement, paired).
Dataset Samples Approximation Time $Q_{1}$ Median MSE $Q_{3}$ MNIST 500 Matrix-Vector $7.140\times 10^{-1}$ $2.625\times 10^{4}$ $2.668\times 10^{4}$ $2.681\times 10^{4}$ Least Squares $1.850\times 10^{0}$ $6.087\times 10^{-2}$ $6.144\times 10^{-2}$ $6.188\times 10^{-2}$ KernelSHAP $7.390\times 10^{-1}$ $9.235\times 10^{1}$ $1.309\times 10^{2}$ $2.147\times 10^{2}$ 1000 Matrix-Vector $9.270\times 10^{-1}$ $8.394\times 10^{3}$ $8.491\times 10^{3}$ $8.566\times 10^{3}$ Least Squares $1.960\times 10^{0}$ $5.879\times 10^{-2}$ $5.914\times 10^{-2}$ $5.949\times 10^{-2}$ KernelSHAP $8.920\times 10^{-1}$ $1.745\times 10^{0}$ $1.784\times 10^{0}$ $1.816\times 10^{0}$ 10000 Matrix-Vector $5.540\times 10^{0}$ $2.385\times 10^{2}$ $2.393\times 10^{2}$ $2.414\times 10^{2}$ Least Squares $7.710\times 10^{0}$ $4.889\times 10^{-2}$ $4.939\times 10^{-2}$ $4.964\times 10^{-2}$ KernelSHAP $2.940\times 10^{0}$ $7.866\times 10^{-2}$ $7.979\times 10^{-2}$ $8.058\times 10^{-2}$ 100000 Matrix-Vector $5.240\times 10^{1}$ $1.048\times 10^{1}$ $1.058\times 10^{1}$ $1.063\times 10^{1}$ Least Squares $6.790\times 10^{1}$ $3.071\times 10^{-2}$ $3.117\times 10^{-2}$ $3.150\times 10^{-2}$ KernelSHAP $2.840\times 10^{1}$ $7.374\times 10^{-3}$ $7.472\times 10^{-3}$ $7.593\times 10^{-3}$ CIFAR10 500 Matrix-Vector $1.990\times 10^{1}$ $1.411\times 10^{3}$ $1.476\times 10^{3}$ $3.191\times 10^{3}$ Least Squares $9.810\times 10^{2}$ $1.440\times 10^{0}$ $1.478\times 10^{0}$ $1.494\times 10^{0}$ KernelSHAP $1.390\times 10^{2}$ $3.112\times 10^{4}$ $1.114\times 10^{5}$ $6.247\times 10^{5}$ 1000 Matrix-Vector $2.180\times 10^{1}$ $5.755\times 10^{2}$ $9.362\times 10^{2}$ $1.314\times 10^{3}$ Least Squares $1.270\times 10^{3}$ $2.025\times 10^{0}$ $2.047\times 10^{0}$ $2.094\times 10^{0}$ KernelSHAP $1.350\times 10^{2}$ $2.411\times 10^{4}$ $1.153\times 10^{5}$ $1.033\times 10^{6}$ 10000 Matrix-Vector $5.690\times 10^{1}$ $9.168\times 10^{1}$ $031\times 10^{2} 1.1$$172\times 10^{2}$ Least Squares $1.210\times 10^{3}$ $1.000\times 10^{1}$ $1.035\times 10^{1}$ $1.059\times 10^{1}$ KernelSHAP $2.820\times 10^{2}$ $2.846\times 10^{1}$ $2.870\times 10^{1}$ $2.897\times 10^{1}$ 100000 Matrix-Vector $3.190\times 10^{2}$ $1.014\times 10^{1}$ $068\times 10^{1} 1.1$$104\times 10^{1}$ Least Squares $3.030\times 10^{3}$ $4.168\times 10^{-1}$ $4.252\times 10^{-1}$ $4.295\times 10^{-1}$ KernelSHAP $1.860\times 10^{3}$ $1.050\times 10^{0}$ $1.053\times 10^{0}$ $1.053\times 10^{0}$
Table 6: Performance Metrics for MNIST and CIFAR-10 Datasets Using Different Methods ($\ell_{2}$-squared estimator, without replacement, paired sampling with $\bm{b}_{0}$).
@@
MNIST 500 Matrix-Vector 0.758 0.721 0.632
Least Squares 0.758 0.718 0.737
KernelSHAP 0.601 0.589 0.959
1000 Matrix-Vector 0.758 0.719 0.639
Least Squares 0.761 0.718 0.742
KernelSHAP 0.707 0.649 0.975
10000 Matrix-Vector 0.762 0.723 0.655
Least Squares 0.762 0.717 0.753
KernelSHAP 0.762 0.710 0.992
100000 Matrix-Vector 0.756 0.715 0.674
Least Squares 0.758 0.719 0.767
KernelSHAP 0.762 0.714 0.996
CIFAR10 500 Matrix-Vector 0.581 0.548 0.008
Least Squares 0.604 0.535 0.033
KernelSHAP 0.607 0.610 0.026
1000 Matrix-Vector 0.560 0.523 0.020
Least Squares 0.584 0.541 0.061
KernelSHAP 0.613 0.612 0.028
10000 Matrix-Vector 0.544 0.510 0.069
Least Squares 0.661 0.571 0.185
KernelSHAP 0.629 0.555 0.184
100000 Matrix-Vector 0.521 0.485 0.239
Least Squares 0.679 0.558 0.662
KernelSHAP 0.685 0.557 0.540
Table 7: AUC and Rank Correlation for MNIST and CIFAR10 Datasets Using Different Methods ($\ell_{2}$-squared estimator, without replacement, paired sampling with $\bm{b}_{0}$).